In a recent year, the total scores for a certain standardized test were normally distributed, with a mean of 500 and a standard deviation of 10.6. Answer below 


(a) Find the probability that a randomly selected medical student who took the test had a total score that was less than 491 The probability that a randomly selected medical student who took the test had a total score that was less than 491 is 0.1977 (Round to four decimal places as needed. ) 

(b) Find the probability that a randomly selected medical student who took the test had a total score that was between 500 and 514 The probability that a randomly selected medical student who took the test had a total score that was between 500 and 514 is 04066 (Round to four decimal places as needed.) 

(c) Find the probability that a randomly selected medical student who took the test had a total score that was more than 522 The probability that a randomly selected medical student who took the test had a total score that was more than 522 is (Round to four decimal places as needed.) Enter your answer in the answer box and then click Check Answer.
Answer #1

mean = = 500

standard deviation = = 10.6

P(x > 522) = 1 - P(x < 522)

= 1 - P((x - ) / < (522 - 500) / 10.6)

= 1 - P(z < 2.08)

= 1 - 0.9812 Using standard normal table.

= 0.0188

Probability = 0.0188